# Explainable Benchmarking for Iterative Optimization Heuristics

### 31 Jan 2024 v1

@@ 3.3. Robustness to Instance Variation and Random Seed Variation
In addition to analyzing the impact of algorithmic components, the approach described in Section 3.2 can be used to study the robustness of an optimization algorithm. Since we collect performance data for multiple runs of each algorithm configuration, the seed used for these runs can be added in as a virtual module of IOH-Xplainer for the analysis. If the algorithm’s performance is independent of the used seed, as is generally expected from the state-of-the-art optimizers, the seed value should have little explanatory power, and thus be assigned a low SHAP value. As such, the seed-based variance  [^2] In IOH-Xplainer, such seed-based variance will be referred to as ‘stochastic variance’ (see Section 4.3.1). is a useful sanity check for algorithm designers.
In addition to the stochasticity inherent to the optimization algorithm, such algorithms are generally run on multiple instances of a problem. These instances can be created, e.g., by rotation and translation of the search space, resulting in slightly different problems, to which an algorithm might or might not be invariant (hansen2011impacts, ; caraffini2019study, ). Thus, looking at the explanatory power of the instance ID (’instance variance’) allows us to check these algorithmic invariances. It should be noted that this analysis is only applicable to benchmark sets which support this type of instance creation, and can thus not directly be used for arbitrary optimization problems.
@@ 4. Explainable Benchmarking of 52 128 Algorithm Configurations
This section discusses benchmarking a large set of heuristic optimisation algorithms by applying the IOH-Xplainer toolbox.
@@ 4.1. Used Algorithms
Table 1. Modules used for modCMA and their options. For population size, default is $4+\lfloor 3\log(d)\rfloor$. Only combinations where $\mu\leq\lambda$ are used.
@@
{default, equal, $w_{i}=1/2^{i}+1/(\lambda 2^{\lambda})$}
@@
$\frac{\lambda}{2}$, 5, 10, 20}
Table 2. Modules used for modDE and their options. For population size, default is $4+\lfloor 3\log(d)\rfloor$.
@@
{‘rand’, ‘best’, ‘target’}
@@
{none, ‘pbest’, ‘best’, ‘rand’}
@@
To illustrate the functionality of IOH-Xplainer, we perform two large-scale experiments on modular optimization algorithms: modular CMA-ES (modCMA) (modcma, ) and modular DE (modDE) (modDE, ). For both algorithms, we perform a full enumeration of a set of available modules and hyper-parameters (categorical values), and we perform a grid of discritized values for the continuous parameters.
@@ 4.3.1. Contribution of hyperparameters
Figure 2. Hyper-parameter contributions per benchmark function for d=5 for modular CMA-ES. Categorical hyper-parameters are encoded to integer values in alphabetic order where NaN is encoded as $-1$. Boolean options are visualized as false (0) and true (1). To be exact, ‘base_sampler’ is either ‘gaussian’ (0), ‘halton’ (1) or ‘sobol’ (2), ‘mirrored’ is either Off ($-1$), ‘mirrored’ (0), or ‘mirrored pairwise’ (1), ‘weights_option’ is either $1/2^{\lambda}$’ (0), ‘default’ (1) or ‘equal’ (2), ‘local_restart’ is either NaN ($-1$), ‘BIPOP’ (0), or ‘IPOP’ (1), and ‘step_size_adaptation’ is either ‘csa’ (0) or ‘psr’ (1).
Figure 3. Hyper-parameter contributions per benchmark function for d=30 for modular CMA-ES. Categorical hyper-parameters are encoded to integer values in alphabetic order where NaN is encoded as $-1$. To be exact, ‘base_sampler’ is either ‘gaussian’ (0), ‘halton’ (1) or ‘sobol’ (2), ‘mirrored’ is either Off ($-1$), ‘mirrored’ (0), or ‘mirrored pairwise’ (1), ‘weights_option’ is either $1/2^{\lambda}$’ (0), ‘default’ (1) or ‘equal’ (2), ‘local_restart’ is either NaN ($-1$), ‘BIPOP’ (0), or ‘IPOP’ (1), and ‘step_size_adaptation’ is either ‘csa’ (0) or ‘psr’ (1).
For several modules, we see a consistent separation between the two colors, such as the ’covariance’ option, where violet dots consistently have negative SHAP values and yellow dots are positive. This indicates that turning off the covariance adaptation mechanism, thus transforming the algorithm into a version of MA-ES (beyer2017simplify, ), leads to worse anytime performance reflected by AOCC. Similarly, many functions have a clear separation between the two options of the ‘elitist’ module. This separation is not always in the same direction, so even though the contribution of this module is consistent on a per-function level, it is not immediately obvious whether elitism is useful for CMA-ES in general.
@@
In both Figure 2 and 3, it is important to point out the differences in the scale of the x-axis. For the ’easier’ functions, e.g. F2, the normalized AOCC can vary by as much as $0.4$, while for other functions, e.g. F24, the overall differences in performance between configurations are much smaller. While it might be natural to assume that for these harder functions, noise effects matter more than module choices, the relative contribution from the stochastic variance (i.e. the seed-based variance) is still quite small and trends in, e.g., ‘elitist’ and ‘weight’ options are rather consistent. We can thus conclude that even for cases where the total performance variation is low, our methodology can still clearly identify parameter contributions.
@@ 4.3.2. Algorithm/configuration selection
f1 Sphere 0.98 (0.00) 0.97 (0.00) 0.74 (0.30) 0.95 (0.00) 0.90 (0.00) 0.66 (0.21)
f2 Ellipsoid 0.91 (0.01) 0.90 (0.00) 0.47 (0.37) 0.29 (0.01) 0.29 (0.01) 0.19 (0.07)
f3 Rastrigin 0.47 (0.41) 0.17 (0.01) 0.12 (0.09) 0.40 (0.01) 0.39 (0.01) 0.35 (0.03)
f4 BuecheRastrigin 0.29 (0.34) 0.13 (0.01) 0.09 (0.04) 0.38 (0.01) 0.38 (0.01) 0.34 (0.02)
f5 LinearSlope 1.00 (0.00) 0.99 (0.00) 0.96 (0.14) 0.99 (0.00) 0.98 (0.00) 0.90 (0.17)
f6 AttractiveSector 0.95 (0.01) 0.91 (0.00) 0.54 (0.36) 0.58 (0.01) 0.58 (0.03) 0.36 (0.11)
f7 StepEllipsoid 0.96 (0.00) 0.95 (0.02) 0.45 (0.33) 0.45 (0.01) 0.45 (0.01) 0.38 (0.04)
f8 Rosenbrock 0.92 (0.01) 0.86 (0.01) 0.49 (0.36) 0.41 (0.01) 0.40 (0.01) 0.33 (0.08)
f9 RosenbrockRotated 0.92 (0.01) 0.87 (0.01) 0.52 (0.34) 0.41 (0.00) 0.41 (0.00) 0.35 (0.07)
f10 EllipsoidRotated 0.91 (0.01) 0.90 (0.01) 0.47 (0.37) 0.29 (0.01) 0.28 (0.01) 0.20 (0.06)
f11 Discus 0.92 (0.01) 0.92 (0.00) 0.44 (0.34) 0.41 (0.01) 0.40 (0.01) 0.35 (0.03)
f12 BentCigar 0.87 (0.03) 0.73 (0.25) 0.41 (0.34) 0.51 (0.09) 0.45 (0.06) 0.24 (0.17)
f13 SharpRidge 0.91 (0.01) 0.88 (0.01) 0.44 (0.33) 0.51 (0.05) 0.49 (0.05) 0.39 (0.07)
f14 DifferentPowers 0.95 (0.01) 0.95 (0.00) 0.67 (0.28) 0.70 (0.01) 0.69 (0.00) 0.56 (0.10)
f15 RastriginRotated 0.49 (0.31) 0.49 (0.31) 0.13 (0.10) 0.40 (0.01) 0.39 (0.01) 0.35 (0.03)
f16 Weierstrass 0.87 (0.09) 0.72 (0.21) 0.26 (0.21) 0.50 (0.02) 0.50 (0.01) 0.43 (0.03)
f17 Schaffers10 0.90 (0.00) 0.90 (0.03) 0.43 (0.24) 0.64 (0.03) 0.60 (0.03) 0.48 (0.05)
f18 Schaffers1000 0.85 (0.00) 0.84 (0.09) 0.31 (0.18) 0.55 (0.02) 0.52 (0.02) 0.44 (0.04)
f19 GriewankRosenbrock 0.41 (0.22) 0.30 (0.02) 0.24 (0.06) 0.55 (0.00) 0.48 (0.01) 0.46 (0.02)
f20 Schwefel 0.46 (0.32) 0.19 (0.01) 0.18 (0.05) 0.48 (0.00) 0.47 (0.00) 0.40 (0.10)
f21 Gallagher101 0.95 (0.05) 0.44 (0.32) 0.36 (0.32) 0.66 (0.22) 0.56 (0.20) 0.45 (0.11)
f22 Gallagher21 0.81 (0.17) 0.62 (0.38) 0.32 (0.30) 0.49 (0.08) 0.44 (0.01) 0.42 (0.03)
f23 Katsuura 0.67 (0.22) 0.48 (0.11) 0.21 (0.11) 0.55 (0.01) 0.49 (0.01) 0.47 (0.03)
f24 LunacekBiRastrigin 0.14 (0.04) 0.12 (0.02) 0.09 (0.03) 0.38 (0.01) 0.38 (0.00) 0.35 (0.02)
Table 4. Performance summary of modular CMA-ES. The gain avg-best is the average improvement of the average-best configuration over all configurations. The gain single-best is the average improvement of the single-best configurations (per function) over all configurations. Boldface in this table indicates a significant improvement (avg-best versus all, and single-best versus avg-best).
@@
Average performance 0.39 0.41
Gain avg-best 0.29 0.09
Gain single-best 0.38 0.11
As observed from the different patterns in Figures 2 and 3, modules show different contribution on different functions. While this means that there is no perfect setting of the modular CMA-ES, it also indicates that an algorithm selector has a lot of potential. To investigate just how large this potential is, we identify the best configuration for each BBOB function (‘single-best’), and the best across all functions (‘avg-best’). The performance for both of these, as well as the average performance over all configurations (‘all’) is shown in Table 3. In Table 4, these performances are aggregated to show the improvement gained by selecting the overall best configuration over the whole set, as well as the improvement which could be gained from a perfect algorithm selector. This table indicates that using the avg-best in modCMA rather than an arbitrary combination would lead to an AOCC improvement of between $0.09$ and $0.29$ for 30 and 5 dimensional BBOB, respectively. With a perfect algorithm selector (operating on settings from Table 1 ), this performance would again be improved, with $0.09$ for 5d and $0.02$ for the 30d setting.
Table 5. Hall of fame of modular CMA-ES: best configurations per function including average best (denoted by All in fid), with resulting values of AOCC and structural BIAS indicator. Denoted in brackets after each module setting is the average effect in AOCC score of choosing the respective option versus all other choices (keeping the rest of the settings the same)
@@
5 All True (0.05) halton (0.01) True (0.11) False (0.29) 20.0 (0.42) IPOP (0.03) mirrored (0.02) 10.0 (0.26) csa (0.06) default (0.08) 0.676553 centre
5 1 False (0.00) halton (0.00) True (0.84) True (0.01) 5.0 (0.05) BIPOP (0.00) mirrored (0.00) 5.0 (0.00) csa (0.02) $1/2^{\lambda}$ (0.45) 0.984495 gaps/clusters
5 2 True (0.03) sobol (0.02) False (0.01) False (0.91) 8.0 (0.11) IPOP (0.00) mirrored (0.02) 5.0 (0.00) csa (0.03) $1/2^{\lambda}$ (0.46) 0.907190 gaps/clusters
5 3 True (0.33) halton (0.29) False (0.31) False (0.43) 8.0 (0.00) BIPOP (0.00) mirrored (0.35) 3.0 (0.23) csa (0.37) default (0.34) 0.472781 centre
5 4 False (0.17) halton (0.17) True (0.26) True (0.16) 8.0 (0.00) IPOP (0.02) pairwise (0.17) 3.0 (0.11) csa (0.19) $1/2^{\lambda}$ (0.17) 0.287226 centre
5 5 True (0.00) gaussian (0.00) True (0.01) False (0.00) 5.0 (0.01) IPOP (0.00) mirrored (0.00) 5.0 (0.00) psr (0.00) default (0.03) 0.998051 -
5 6 True (0.01) halton (0.02) True (0.91) True (0.18) 5.0 (0.12) BIPOP (0.00) pairwise (0.89) 5.0 (0.01) csa (0.16) default (0.69) 0.953062 gaps/clusters
5 7 True (0.15) halton (0.37) False (0.00) False (0.77) 20.0 (0.68) BIPOP (0.00) mirrored (0.28) 10.0 (0.43) psr (0.28) default (0.22) 0.962005 -
5 8 True (0.07) halton (0.09) True (0.92) True (0.09) 8.0 (0.28) BIPOP (0.00) pairwise (0.11) 5.0 (0.07) csa (0.36) default (0.42) 0.922699 gaps/clusters
5 9 False (0.02) sobol (0.03) False (0.06) False (0.84) 5.0 (0.00) IPOP (0.03) mirrored (0.03) 2.0 (0.72) csa (0.11) default (0.06) 0.915234 gaps/clusters
5 10 False (0.00) halton (0.01) True (0.02) True (0.91) 5.0 (0.17) BIPOP (0.00) mirrored (0.01) 5.0 (0.08) psr (0.06) equal (0.47) 0.906337 -
5 11 True (0.10) sobol (0.01) True (0.02) False (0.63) 20.0 (0.77) IPOP (0.00) mirrored (0.05) 10.0 (0.47) csa (0.02) default (0.05) 0.917868 centre
5 12 True (0.06) halton (0.04) False (0.02) False (0.87) 8.0 (0.00) BIPOP (0.00) mirrored (0.08) 3.0 (0.02) csa (0.13) default (0.16) 0.870508 centre
5 13 False (0.00) halton (0.02) True (0.02) True (0.91) 5.0 (0.20) BIPOP (0.00) mirrored (0.02) 5.0 (0.32) psr (0.14) equal (0.48) 0.905252 -
5 14 False (0.30) sobol (0.01) True (0.77) True (0.78) 5.0 (0.07) IPOP (0.00) pairwise (0.01) 5.0 (0.02) csa (0.01) equal (0.03) 0.954450 centre
5 15 True (0.29) halton (0.10) True (0.33) False (0.35) 20.0 (0.35) IPOP (0.12) mirrored (0.15) 10.0 (0.31) csa (0.32) default (0.27) 0.492284 centre
5 16 True (0.13) halton (0.14) True (0.06) False (0.29) 20.0 (0.44) BIPOP (0.04) mirrored (0.13) 5.0 (0.31) csa (0.49) default (0.42) 0.866222 centre
5 17 False (0.01) halton (0.02) True (0.05) False (0.11) 20.0 (0.51) BIPOP (0.00) mirrored (0.03) 10.0 (0.19) csa (0.07) default (0.13) 0.904906 centre
5 18 True (0.15) halton (0.14) True (0.22) False (0.41) 20.0 (0.56) BIPOP (0.00) mirrored (0.31) 5.0 (0.21) psr (0.09) default (0.45) 0.850612 -
5 19 False (0.16) halton (0.15) True (0.16) False (0.10) 20.0 (0.10) BIPOP (0.03) mirrored (0.16) 5.0 (0.16) csa (0.17) $1/2^{\lambda}$ (0.11) 0.412441 discretisation
5 20 False (0.24) halton (0.26) True (0.30) True (0.27) 8.0 (0.22) IPOP (0.15) pairwise (0.26) 5.0 (0.16) csa (0.27) default (0.28) 0.463041 centre
5 21 False (0.21) halton (0.31) True (0.84) True (0.26) 5.0 (0.00) IPOP (0.16) mirrored (0.32) 2.0 (0.84) csa (0.66) default (0.30) 0.951148 centre
5 22 True (0.11) halton (0.14) True (0.75) True (0.46) 5.0 (0.00) IPOP (0.21) pairwise (0.30) 2.0 (0.16) csa (0.52) $1/2^{\lambda}$ (0.21) 0.812444 centre
5 23 False (0.23) halton (0.37) False (0.22) False (0.51) 20.0 (0.44) BIPOP (0.00) mirrored (0.32) 5.0 (0.28) psr (0.44) default (0.32) 0.674847 -
5 24 False (0.00) halton (0.02) True (0.01) False (0.03) 8.0 (0.00) IPOP (0.02) mirrored (0.01) 3.0 (0.02) csa (0.04) $1/2^{\lambda}$ (0.02) 0.141660 gaps/clusters
30 All True (0.01) sobol (0.05) False (0.00) False (0.16) 14.0 (0.00) IPOP (0.00) mirrored (0.01) 4.0 (0.01) csa (0.01) default (0.01) 0.496598 gaps/clusters
30 1 False (0.00) sobol (0.12) False (0.01) False (0.06) 5.0 (0.11) IPOP (0.00) mirrored (0.01) 5.0 (0.01) psr (0.01) $1/2^{\lambda}$ (0.34) 0.945324 -
30 2 True (0.00) sobol (0.06) False (0.01) True (0.22) 10.0 (0.07) IPOP (0.00) pairwise (0.01) 10.0 (0.04) psr (0.21) equal (0.10) 0.294408 -
30 3 False (0.00) sobol (0.04) False (0.00) False (0.07) 20.0 (0.04) BIPOP (0.00) mirrored (0.01) 10.0 (0.02) csa (0.00) default (0.01) 0.396627 centre
30 4 False (0.00) sobol (0.04) False (0.00) False (0.06) 20.0 (0.03) BIPOP (0.00) mirrored (0.00) 10.0 (0.01) csa (0.01) default (0.01) 0.382659 gaps/clusters
30 5 False (0.06) sobol (0.02) True (0.06) True (0.01) 5.0 (0.03) IPOP (0.00) mirrored (0.01) 5.0 (0.01) psr (0.58) default (0.01) 0.992649 -
30 6 True (0.02) sobol (0.14) False (0.03) False (0.38) 14.0 (0.00) BIPOP (0.00) mirrored (0.03) 4.0 (0.01) csa (0.10) default (0.07) 0.583744 gaps/clusters
30 7 True (0.00) sobol (0.04) False (0.01) False (0.12) 14.0 (0.02) IPOP (0.01) pairwise (0.00) 5.0 (0.00) csa (0.04) default (0.01) 0.447991 gaps/clusters
30 8 True (0.01) sobol (0.01) False (0.01) False (0.22) 5.0 (0.02) IPOP (0.01) mirrored (0.01) 5.0 (0.01) csa (0.02) $1/2^{\lambda}$ (0.13) 0.409491 discretisation
30 9 False (0.01) sobol (0.02) True (0.16) True (0.06) 5.0 (0.01) IPOP (0.01) mirrored (0.00) 5.0 (0.01) csa (0.00) equal (0.07) 0.409810 gaps/clusters
30 10 True (0.00) sobol (0.03) False (0.01) True (0.20) 10.0 (0.06) IPOP (0.00) pairwise (0.00) 10.0 (0.03) psr (0.19) equal (0.09) 0.289573 -
30 11 True (0.01) sobol (0.03) True (0.04) False (0.05) 14.0 (0.00) BIPOP (0.00) mirrored (0.02) 4.0 (0.00) psr (0.01) default (0.01) 0.408763 -
30 12 False (0.03) sobol (0.11) True (0.51) True (0.04) 5.0 (0.13) BIPOP (0.05) mirrored (0.05) 5.0 (0.05) csa (0.11) $1/2^{\lambda}$ (0.22) 0.505465 gaps/clusters
30 13 True (0.03) sobol (0.06) False (0.04) False (0.07) 14.0 (0.00) BIPOP (0.03) mirrored (0.04) 4.0 (0.02) psr (0.03) default (0.05) 0.511117 -
30 14 False (0.00) gaussian (0.03) True (0.31) True (0.02) 5.0 (0.06) BIPOP (0.00) mirrored (0.01) 5.0 (0.01) csa (0.03) $1/2^{\lambda}$ (0.14) 0.704631 gaps/clusters
30 15 True (0.00) sobol (0.01) False (0.01) False (0.03) 20.0 (0.04) IPOP (0.00) mirrored (0.01) 10.0 (0.02) psr (0.00) default (0.01) 0.398682 -
30 16 True (0.00) sobol (0.03) False (0.01) False (0.10) 20.0 (0.06) IPOP (0.00) mirrored (0.02) 10.0 (0.02) csa (0.01) default (0.03) 0.499925 centre
30 17 True (0.00) sobol (0.08) False (0.02) False (0.20) 20.0 (0.11) BIPOP (0.00) mirrored (0.01) 10.0 (0.03) csa (0.02) default (0.06) 0.635100 centre
30 18 True (0.00) sobol (0.07) False (0.00) False (0.15) 20.0 (0.09) BIPOP (0.00) pairwise (0.01) 10.0 (0.03) csa (0.03) default (0.02) 0.549818 gaps/clusters
30 19 True (0.00) gaussian (0.01) True (0.10) True (0.03) 5.0 (0.03) BIPOP (0.09) mirrored (0.01) 5.0 (0.00) csa (0.10) equal (0.04) 0.548607 gaps/clusters
30 20 True (0.00) sobol (0.01) True (0.26) True (0.00) 5.0 (0.01) IPOP (0.00) mirrored (0.00) 5.0 (0.00) csa (0.00) $1/2^{\lambda}$ (0.04) 0.481448 gaps/clusters
30 21 True (0.12) sobol (0.17) True (0.28) True (0.18) 14.0 (0.00) IPOP (0.19) mirrored (0.15) 4.0 (0.09) csa (0.15) $1/2^{\lambda}$ (0.14) 0.664094 gaps/clusters
@@
30 22 False (0.05) halton (0.04) True (0.05) False (0.05) 10.0 (0.05) BIPOP (0.00) mirrored (0.06) 5.0 (0.05) psr (0.06) $1/2^{\lambda}$ (0.05) 0.485074 -
30 23 False (0.01) sobol (0.01) True (0.00) False (0.09) 20.0 (0.05) IPOP (0.00) pairwise (0.01) 10.0 (0.01) psr (0.09) default (0.01) 0.549599 -

### 23 Feb 2024 v2

@@ 3.3. Robustness to Instance Variation and Random Seed Variation
In addition to analyzing the impact of algorithmic components, the approach described in Section 3.2 can be used to study the robustness of an optimization algorithm. Since we collect performance data for multiple runs of each algorithm configuration, the seed used for these runs can be added in as a virtual module of IOHxplainer for the analysis. If the algorithm’s performance is independent of the used seed, as is generally expected from the state-of-the-art optimizers, the seed value should have little explanatory power and thus be assigned a low SHAP value. As such, the seed-based variance  [^1] In IOHxplainer, such seed-based variance will be referred to as ‘stochastic variance’ (see Section 4.3.1). is a useful sanity check for algorithm designers.
In addition to the stochasticity inherent to the optimization algorithm, such algorithms are generally run on multiple instances of a problem. These instances can be created, e.g., by rotation and translation of the search space, resulting in slightly different problems, to which an algorithm might or might not be invariant (hansen2011impacts, ; caraffini2019study, ). Thus, looking at the explanatory power of the instance ID (‘instance variance’) allows us to check these algorithmic invariances. It should be noted that this analysis is only applicable to benchmark sets which support this type of instance creation and can thus not directly be used for arbitrary optimization problems.
@@ 4. Explainable Benchmarking of 52 128 Algorithm Configurations
This section discusses benchmarking a large set of heuristic optimisation algorithms by applying the IOHxplainer toolbox.
@@ 4.1. Used Algorithms
Table 1. Modules used for modCMA and their options. Only combinations where $\mu\leq\lambda$ are used. Values of default populations sizes are computed according to $4+\lfloor 3\log(d)\rfloor$ for dimensionalities $d\in\{5,30\}$ and marked as ${}^{\star}$, while resulting offspring sizes are computed as $\frac{\lambda}{2}$ and marked with ${}^{\dagger}$. Values of recombination weights for the $\lambda$-decay are computed as $w_{i}=\frac{1}{2^{i}}+\frac{1}{\lambda 2^{\lambda}}$. Colour coding serves to distinguish options within each module in Figures 2, 3 and Table 5.
@@
{default, equal, $\lambda$-decay}
@@
{${}^{\dagger}$, ${}^{\dagger}$, ${}^{\dagger}$, ${}^{\dagger}$, ${}^{\dagger}$, 20, ${}^{\dagger}$}
Table 2. Modules used for modDE and their options. Values of default populations sizes are computed according to $4+\lfloor 3\log(d)\rfloor$ for dimensionalities $d\in\{5,30\}$ and marked as ${}^{\star}$. Colour coding serves to distinguish options within each module in Figures 4, 5 and Table 8.
@@
{best, rand, target}
@@
{none, best, pbest, rand}
@@
To illustrate the functionality of IOHxplainer, we perform two large-scale experiments on modular optimization algorithms: modular CMA-ES (modCMA) (modcma, ) and modular DE (modDE) (modDE, ). For both algorithms, we perform a full enumeration of a set of available modules and hyperparameters (categorical values) and we perform a grid of discritized values for the continuous parameters.
@@ 4.3.1. Contribution of hyperparameters
Figure 2. Hyper-parameter contributions per benchmark function for d=5 for modular CMA-ES. Categorical hyperparameters are encoded to integer values in alphabetic order where NaN is encoded as $-1$. Refer to Table 2 for the colour coding used.
Figure 3. Hyper-parameter contributions per benchmark function for d=30 for modular CMA-ES. Categorical hyperparameters are encoded to integer values in alphabetic order where NaN is encoded as $-1$. Refer to Table 2 for the colour coding used.
For several modules, we see a consistent separation between the two colors, such as the ‘covariance’ option, where violet dots consistently have negative SHAP values and yellow dots are positive. This indicates that turning off the covariance adaptation mechanism, thus transforming the algorithm into a version of MA-ES (beyer2017simplify, ), leads to worse anytime performance reflected by AOCC. Similarly, many functions have a clear separation between the two options of the ‘elitist’ module. This separation is not always in the same direction, so even though the contribution of this module is consistent on a per-function level, it is not immediately obvious whether elitism is useful for CMA-ES in general.
@@
In both Figure 2 and 3, it is important to point out the differences in the scale of the x-axis. For the ‘easier’ functions, e.g. F2, the normalized AOCC can vary by as much as $0.4$, while for other functions, e.g. F24, the overall differences in performance between configurations are much smaller. While it might be natural to assume that for these harder functions, noise effects matter more than module choices, the relative contribution from the stochastic variance (i.e. the seed-based variance) is still quite small and trends in, e.g., ‘elitist’ and ‘weight’ options are rather consistent. We can thus conclude that even for cases where the total performance variation is low, our methodology can still clearly identify parameter contributions.
@@ 4.3.2. Algorithm/configuration selection
f1 Sphere 0.98 (0.00) 0.97 (0.00) 0.69 (0.30) 0.95 (0.00) 0.90 (0.00) 0.62 (0.21)
f2 Ellipsoid 0.91 (0.01) 0.90 (0.01) 0.39 (0.36) 0.29 (0.01) 0.29 (0.01) 0.18 (0.06)
f3 Rastrigin 0.38 (0.29) 0.30 (0.23) 0.12 (0.08) 0.40 (0.00) 0.39 (0.01) 0.35 (0.02)
f4 BuecheRastrigin 0.17 (0.16) 0.13 (0.02) 0.09 (0.04) 0.38 (0.01) 0.38 (0.01) 0.34 (0.02)
f5 LinearSlope 1.00 (0.00) 0.99 (0.00) 0.95 (0.16) 0.99 (0.00) 0.98 (0.00) 0.89 (0.17)
f6 AttractiveSector 0.95 (0.01) 0.90 (0.01) 0.48 (0.34) 0.58 (0.01) 0.58 (0.03) 0.35 (0.11)
f7 StepEllipsoid 0.95 (0.00) 0.94 (0.03) 0.50 (0.31) 0.45 (0.01) 0.45 (0.01) 0.38 (0.04)
f8 Rosenbrock 0.91 (0.02) 0.85 (0.02) 0.42 (0.34) 0.41 (0.01) 0.40 (0.01) 0.32 (0.08)
f9 RosenbrockRotated 0.92 (0.01) 0.86 (0.01) 0.45 (0.33) 0.41 (0.00) 0.41 (0.00) 0.34 (0.06)
f10 EllipsoidRotated 0.90 (0.01) 0.89 (0.01) 0.39 (0.36) 0.29 (0.01) 0.28 (0.01) 0.19 (0.06)
f11 Discus 0.92 (0.01) 0.91 (0.01) 0.37 (0.32) 0.41 (0.01) 0.40 (0.01) 0.35 (0.02)
f12 BentCigar 0.86 (0.07) 0.73 (0.23) 0.33 (0.33) 0.51 (0.09) 0.45 (0.06) 0.21 (0.17)
f13 SharpRidge 0.90 (0.01) 0.87 (0.01) 0.37 (0.31) 0.51 (0.05) 0.49 (0.05) 0.38 (0.07)
f14 DifferentPowers 0.95 (0.01) 0.95 (0.00) 0.62 (0.27) 0.70 (0.01) 0.69 (0.00) 0.54 (0.10)
f15 RastriginRotated 0.48 (0.02) 0.43 (0.19) 0.13 (0.10) 0.40 (0.01) 0.39 (0.01) 0.35 (0.02)
f16 Weierstrass 0.81 (0.16) 0.69 (0.19) 0.26 (0.19) 0.50 (0.01) 0.50 (0.01) 0.43 (0.03)
f17 Schaffers10 0.90 (0.02) 0.87 (0.05) 0.41 (0.21) 0.64 (0.02) 0.60 (0.03) 0.48 (0.04)
f18 Schaffers1000 0.82 (0.12) 0.68 (0.16) 0.30 (0.16) 0.55 (0.01) 0.52 (0.02) 0.44 (0.04)
f19 GriewankRosenbrock 0.40 (0.23) 0.31 (0.02) 0.24 (0.06) 0.55 (0.00) 0.48 (0.01) 0.46 (0.02)
f20 Schwefel 0.36 (0.11) 0.21 (0.07) 0.19 (0.05) 0.48 (0.00) 0.47 (0.00) 0.40 (0.09)
f21 Gallagher101 0.84 (0.23) 0.53 (0.32) 0.40 (0.31) 0.66 (0.22) 0.56 (0.20) 0.45 (0.10)
f22 Gallagher21 0.79 (0.26) 0.46 (0.33) 0.36 (0.30) 0.48 (0.11) 0.44 (0.01) 0.42 (0.03)
f23 Katsuura 0.66 (0.22) 0.49 (0.20) 0.21 (0.10) 0.55 (0.01) 0.49 (0.01) 0.47 (0.03)
f24 LunacekBiRastrigin 0.14 (0.03) 0.13 (0.02) 0.09 (0.02) 0.38 (0.00) 0.38 (0.00) 0.35 (0.01)
Table 4. Performance summary of modular CMA-ES. The gain avg-best is the average improvement of the average-best configuration over all configurations. The gain single-best is the average improvement of the single-best configurations (per function) over all configurations. Boldface in this table indicates a significant improvement (avg-best versus all and single-best versus avg-best).
@@
Average performance 0.37 0.40
Gain avg-best 0.30 0.09
Gain single-best 0.38 0.12
As observed from the different patterns in Figures 2 and 3, modules show different contribution on different functions. While this means that there is no perfect setting of the modular CMA-ES, it also indicates that an algorithm selector has a lot of potential. To investigate just how large this potential is, we identify the best configuration for each BBOB function (‘single-best’) and the best across all functions (‘avg-best’). The performance for both of these, as well as the average performance over all configurations (‘all’) is shown in Table 3. In Table 4, these performances are aggregated to show the improvement gained by selecting the overall best configuration over the whole set, as well as the improvement which could be gained from a perfect algorithm selector. This table indicates that using the avg-best in modCMA rather than an arbitrary combination would lead to an AOCC improvement of between $0.09$ and $0.30$ for 30 and 5 dimensional BBOB, respectively. With a perfect algorithm selector (operating on settings from Table 1 ), this performance would again be improved, with $0.09$ for 5d and $0.03$ for the 30d setting.
Table 5. Hall of fame of modular CMA-ES: best configurations per function including average best (denoted by All in fid), with resulting values of AOCC and structural BIAS indicator. Denoted in brackets after each module setting is the average effect in AOCC score of choosing the respective option versus all other choices (keeping the rest of the settings the same). Colouring of module options is consistent with Table 1 and Figures 2,  3. Options marked in italic are consistently selected on this function for both dimensionalities.
@@
5 All true (0.04) Gaussian (0.19) true (0.13) false (0.28) 20 (0.38) IPOP (0.02) mirrored (0.05) 10 (0.25) CSA (0.08) default (0.09) 0.666806 centre
@@
5 1 false (0.00) Halton (0.00) true (0.85) true (0.00) 5 (0.10) BIPOP (0.00) mirrored (0.00) 5 (0.00) CSA (0.08) $\lambda$-decay (0.16) 0.983122 centre